Two blocks of mass m1 nd m2 are connected to each other through a light spring. If we push m1 with force f then its acceleration is a1. What will be the acceleration of m2

Grade:12th pass

1 Answers

Susmita
5 years ago
We are applying an external force f on mass m1.To cause an acceleration,we have to overcome the restoring force due to the spring which is kx.So (f-kx) this excess amount of force will cause the acceleration.From the force equation
m1a1=f-kx
or,kx=f-m1a1
As m1 is displaced,it will try to displace m2.But m2 will try to maintain its position due to restoring force kx.So on m2 the only force acting is kx.For m2 the firce equation is
f'=kx
Using the previous equation for m1,replacing the value of kx we get
f'=f-m1a1
So acceleration on m2 is
a2=f'/m2=(f-m1a1)/m2
please approve the answer if you are helped.